# Policy Learning with a Language Bottleneck

### v2

@@ 8.2 PLLB can learn counter-intuitive policies
8.3 PLLB learns more generalizable policies
To evaluate generalization, we create an environment set-up with a new bag, location, and set of visual distractions (see Appendix 15). As shown in Figure 11, the rules learned by Bottleneck outperform alternative methods for generalizing to this new environment. Because the Vanilla RL policy can only rely on image features, it is not able to leverage the more generalizable strategy PLLB rules capture. Overall, the results in Grasp show that PLLB can improve policy convergence even in the complex dynamics of embodied tasks like robotic grasping.
@@ 9 Limitations & Discussion
PLLB helps agents becomes more interpretable by and inter-operable with humans. In our current implementation, rule_gen requires converting agent episodes into representations an LM can handle (e.g., text or images), which may prevent the use of long-horizon sensorimotor trajectories. Nevertheless, our Grasp task shows how PLLB can leverage progress in pretrained models for robotic tasks, and future advances in long context modeling can address this limitation. Rule generation can also fail when the LM abstracts too much (e.g., in Birds we found the rule: “Use descriptive language that conveys the mood, such as the serenity of a snowy day or the freedom of flight”). Advances in multi-modal models and better approaches for selecting contrastive examples can also mitigate these issues (Appendix B.2).
An interesting direction for future work is applying PLLB to tasks with complex reward functions that reflect hard-to-articulate human preferences, such as image captioning for accessibility (Nie et al., 2020) and personalized language models (Li et al., 2023a). Linguistic rules could enhance transparency and predictability even when users struggle to express their goals.
Additionally, PLLB could generate not just abstract rules but also explanations, goals, or learning strategies, enabling cultural transmission or novel uses in update functions—such as mapping language to state/action abstractions, goals, or reward functions. This paves the way for more human-like AI agents that creatively integrate language with decision-making (Colas et al., 2022).
Acknowledgements
@@
Appendix A Additional Methods
A.1 Hyperparameters
@@
A.2 Maze
See Figures 12 and 13.
@@ A.2 Maze
A.3 Dataset Details for Birds.
@@ A.4 Grasp
A.5 Full Prompts
@@
    SaySelect: $[\mathcal{L}]$+Agent 1 selected [observation]. So I should select
@@
    RULE: + $[\mathcal{L}]$+ PREVIOUS ACTIONS: + $[\tau_{1...t-1}]$ + CURRENT OBSERVATION: + [observation] +
@@
    Birds: Provide a one-sentence description of this image, using the following $\mathcal{L}$
@@
Appendix B Complementary Results
B.1 Examples of Generated Rules for All Environments
@@
B.2 Effect of Model on Generated Rules

### v3

@@ 8.2 PLLB can learn counter-intuitive policies
8.3 PLLB improves few-shot generalization
To evaluate few0shot generalization, we create an environment with a new bag, location, and set of visual distractions (see Appendix 12). We then evaluate Bottleneck, Zero-Shot VLM (Oracle), and Vanilla RL (Baseline) when intialized with the policies learned in the original environment, allowing for updates from the environment over time. As shown in Figure 7, the rules learned by Bottleneck outperform alternative methods for generalizing to this new environment. Because the Vanilla RL policy can only rely on image features, it is not able to leverage the more generalizable strategy PLLB rules capture and select grasp keypoints that hold semantic meaning (e.g. the bottom of the bag) captured by pre-trained LMs. Overall, the Grasp results show that PLLB can improve policy convergence even for the complex dynamics of embodied tasks.
@@ 9 Limitations & Discussion
PLLB helps agents becomes more interpretable by and inter-operable with humans. In our current implementation, rule_gen requires converting episodes into LM-compatible representations (text/images), which may limit use with long-horizon sensorimotor trajectories. Nevertheless, our Grasp task successfully uses pretrained robotic models, and advances in long-context modeling can address this limitation. A natural question is whether PLLB’s gains arise specifically from language or from reward-aligned abstraction more generally. Comparing against non-linguistic bottlenecks (e.g., learned latent codes) is challenging: such representations require training across many task variations to become meaningful, whereas language provides pre-structured abstractions via pretrained LLMs that transfer to single tasks. Our Adversarial ablations across experiments confirm that rule content, not merely the presence of a bottleneck, drives improvements with PLLB.
Our human studies also demonstrate uniquely linguistic benefits (interpretability, human transfer) that would be unavailable with other latent representations (including code). However, rule generation can fail when the LM either abstracts too much (e.g., in Birds we once found the rule: “Use descriptive language that conveys the mood, such as the serenity of a snowy day or the freedom of flight”) or, more seriously, causes harm in safety-critical situations by creating a false trust in generated rules. We believe improvements in multi-modal models can help mitigate these issues (Appendix B.9), as well as additional human-in-the-loop verification steps of rules for high-stakes situations.
Our work opens up a set of interesting questions around designing intelligent sampling of contrastive episodes in rule_gen. For example, sampling episodes that are not outliers can help with handling stochastic environments, where high reward trajectories may not necessarily be optimal. Sampling of contrastive episodes can also adjust to known properties of how LMs attend to long context (i.e. more attention to episodes at the start and end). Another interesting direction for future work is applying PLLB to tasks with complex reward functions that reflect hard-to-articulate human preferences, such as image captioning for accessibility (Nie et al., 2020) and personalizing language models (Li et al., 2023a). Although one concern is whether the linguistic rules from PLLB might exacerbate spurious reward signals, any reward misspecification becomes more detectable and diagnosable through interpretable rules, whereas black-box policies offer no such transparency.
Our experiments with PLLB focus on environments with tractable action spaces (e.g. discrete choices, text captions, keypoints for robotic grasping). In like open-ended text generation in math reasoning or code synthesis, this approach would require adaptation. Possible extensions include: (1) using the rule as a verifier or reward bonus rather than an action prior, (2) discretizing the action space into semantically meaningful categories that the LM can reason over, or (3) using the rule to filter or re-rank candidate actions from a base policy. Such adaptations would extend the cognitive and communicative benefits of language-guided learning to a broader class of sequential decision-making problems (Colas et al., 2022), and we leave exploration of these extensions to future work.
10 Acknowledgements
@@
Appendix A Appendix
Appendix B Appendix
B.1 Hyperparameters
@@
B.2 SelectSay Overview
B.3 Maze
See Figures 9 and 10.
@@ B.3 Maze
B.4 Dataset Details for Birds.
@@ B.5 Grasp
B.6 Builder and Birds Iterations
B.7 Full Prompts
@@
    SaySelect: $\mathcal{L}$]+Agent 1 selected [observation]. So I should select
@@
    RULE: + $\mathcal{L}$]+ PREVIOUS ACTIONS: + $\tau_{1...t-1}]$ + CURRENT OBSERVATION: + [observation] +
@@
[…]
@@
B.8 Examples of Generated Rules for All Environments
@@
B.9 Effect of Model on Generated Rules
@@
B.10 Sensitivity Analysis of Rule Generation
We next compare the sensitivity of PLLB to the implementation of gen_rule, specifically variations in the prompt and temperature for sampling. We present results for SaySelect, as its discrete action space and deterministic dynamics provide the clearest signal for detecting sensitivity effects.
B.10.1 Prompt Variations
For SaySelect, we evaluate prompt sensitivity in gen_rule by comparing the following variations with the Adversarial baseline described in 5.:
  • (Original) You will be given a list of (OBSERVATION, ACTION, REWARD) examples collected from two agents learning to solve a task. Possible ACTIONS an agent can take are: 1, 2, 3, 4, 5, and quit. Each OBSERVATION describes the ordered sequence of actions that AGENT 1 picks, and each ACTION describes the ACTION that AGENT 2 picks based on the given OBSERVATION. The examples are separated into HIGH REWARD and LOW REWARD examples.+[samples]+Output a language rule that best summarizes the strategy AGENT 2 should follow to receive HIGH REWARD, not LOW REWARD, based on the examples. Start the instruction with the prefix ‘I should’.
  • (No Format Instruction) You will be given a list of (OBSERVATION, ACTION, REWARD) examples collected from two agents learning to solve a task. Possible ACTIONS an agent can take are: 1, 2, 3, 4, 5, and quit. Each OBSERVATION describes the ordered sequence of actions that AGENT 1 picks, and each ACTION describes the ACTION that AGENT 2 picks based on the given OBSERVATION. The examples are separated into HIGH REWARD and LOW REWARD examples.+[samples]+Output a language rule that best summarizes the strategy AGENT 2 should follow to receive HIGH REWARD, not LOW REWARD, based on the examples.
  • (Low Context) You will be given a list of (OBSERVATION, ACTION, REWARD) examples collected from two agents learning to solve a task. Output a language rule that best summarizes the strategy AGENT 2 should follow to receive HIGH REWARD, not LOW REWARD, based on the examples. Start the instruction with the prefix ‘I should’.
  • (Rephrase) You’ve been given a list of (OBSERVATION, ACTION, REWARD) triples from two agents learning to solve a task. Possible ACTIONS each agent might take are: 1, 2, 3, 4, 5, and quit. Each OBSERVATION refers to the ordered sequence of actions that AGENT 1 selects, and each ACTION refers to the ACTION that AGENT 2 selects based on the seen OBSERVATION. The examples are divided into HIGH REWARD and LOW REWARD examples.+[samples]+Describe the strategy AGENT 2 uses in HIGH REWARD examples that differs from LOW REWARD examples. Start it with ‘I should’.
Table 6 and 7 show that PLLB is mostly robust to variations in prompt and always outperforms the Adversarial baseline, showing that PLLB indeed places higher weight on the content of the contrasting episodes rather than relying on any specific syntax for rule generation.
B.10.2 Temperature Variation
For SaySelect, our default temperature for sampling with gen_rule is 0.5. We compare with temperatures 0.1 (low diversity) and 0.9 (high diversity), and find only slight decrease in reward at the higher temperature of 0.9 (see Table 6). Importantly, across all temperatures the average reward and human-interpretability scores are stronger than baselines, indicating robustness towards sampling temperature.
Episode 500 1000 6000
[…]
Table 6: Reward across ablations to probe sensitivity of rule generation to prompt variation and temperature. Results are averages across 5 trials.
Episode 500 1000 6000
[…]
Interpretability (Original, temp=0.9) $0.72\pm 0.05$ $0.80\pm 0.05$ $0.93\pm 0.03$
[…]
Interpretability (No Format Instruction) $0.67\pm 0.06$ $0.91\pm 0.03$ $0.94\pm 0.02$
[…]
Table 7: Human interpretability across ablations to probe sensitivity of rule generation to prompt variation and temperature. Results are averages across 5 trials.
B.11 Non-Contrastive Variant
A core feature of PLLB is using contrastive episodes (examples of both low and high reward episodes) to gen_rule. Here we report a non-contrastive ablation for the SaySelect (RL) and Birds (diffusion-model based communication game) environments, where we only use high reward examples in the prompt to gen_rule. Our results show that contrastive episodes are important for strong performance. Concretely, for SaySelect, the non-contrastive ablation decreases both task reward and human-interpretability of the RL policy, across different total number of episodes seen (see Table 8). Likewise, for the Birds task, the non-contrastive ablation results in lower reward than PLLB, although still leads to a non-trivial improvement over the Baseline and Adversarial methods (see Table 9). Upon a closer look, we find that without contrasting episodes (i.e. negative episodes that highlight what behaviors are shared between high and low rewarding strategies), PLLB rules often capture spurious features of the tasks, such as “ Include details about the bird’s actions. This helps in creating a more vivid and engaging description.”. This subsequently results in highly-specific captions such as a bird ‘possibly displaying a territorial or mating display‘ or ‘looking at the camera‘, neither of which are relevant for any of our reward functions. Overall, these results highlight the importance of using contrasting episodes to improve PLLB rule equality.
Episode 500 1000 6000
[…]
Reward (Non-contrastive) $0.53\pm 0.02$ $0.7\pm 0.2$ $0.91\pm 0.03$
[…]
Interpretability (Non-contrastive) $0.75\pm 0.02$ $\pm 0.06 0.88±0$$8\pm 0.02$
[…]
Table 8: Reward and human-interpretability scores for contrastive and non-contrastive versions of PLLB in SaySelect. Results are averages across 5 trials.
(a) Color
Method Score
Baseline $1.69\pm 0.5$
[…]
Non-contrastive $2.64\pm 0.3$
[…]
(b) Background
Method Score
Baseline $1.27\pm 0.5$
[…]
Non-contrastive $1.78\pm 0.4$
[…]
(c) Species
Method Score
Baseline $1.59\pm 0.5$
[…]
Non-contrastive $2.51\pm 0.5$
[…]
Table 9: Noncontrastive ablation in comparison with main results for Birds task across 3 reward functions. Results are averaged across 5 trials.